OpenStudy (anonymous):

Find the solution, given the replacement set. 5x - 2y = 6 {(-2, -8), (0, -3), (1, -1), (2, 4), (4, 7)} A. {(-2, -8), (2, 4), (4, 7)} B. {(-2, -8), (0, -3), (4, 7)} C. {(-2, -8), (0, -3), (1, -1)} D. {(1, -1), (2, 4), (4, 7)}

OpenStudy (campbell_st):

substitute each ordered pair and find which gives a value of 6 there is more than 1 ordered pair that is a solution e.g. use (-2, -8) so 5*(-2) -2*(-8) = -10 + 16 = 6 so (-2, -8) is a solution ' try (1, -1) 5*1 -2*(-1) = 5 + 2 = 7' not in the solution set. Hope this helps you find the correct answer